# Gauge potentials and vortices in the Fock space of a pair of periodically driven Bose-Einstein condensates

### v1

@@ IV.1 Eigenenergies and Eigenstates
In this paper we focus on the ground state properties of the system, so we take $n_{r}=0$ from now on. When $u<1$, the ground state energy is $E_{0,0}=\omega$ and the wave function is simply a Gaussian centered at $r=0$. In terms of the two BECs, it is centered at $m=n=0$ which is the state with an equal number of bosons of each BEC in each well. When $u>1$, the centrifugal limit is surpassed and Eq. (12) predicts that $E_{0,m_{z}}\to-\infty$ as $m_{z}\to\infty$. However, this is only true in the thermodynamic limit. For finite sized BECs there can be a unique ground state with $m_{z}\neq 0$ which, according to Eq. (13), is a vortex state with $m_{z}$ quanta of angular momentum.
@@
$\hat{L}^{\mathrm{MB}}_{z}$ represents the angular momentum operator in the FSL and we expect its mean value $\langle\hat{L}^{\mathrm{MB}}_{z}\rangle$ to accurately give $m_{z}$ when computed for the low lying energy states well away from the edge of the FSL, which are the states of primary interest in this paper. However, we note that the correspondence between $\langle\hat{L}^{\mathrm{MB}}_{z}\rangle$ and $m_{z}$ can break down for higher states due to the omission of terms that depend on the number of particles as $1/N,1/N^{3/2},\ldots$. These terms were dropped in the derivation of $\hat{H}_{\mathrm{eff}}$ and lead to corrections that depend on the quadrature variables as $x^{2}y^{2},x^{2}p_{y}^{2},y^{2}p_{x}^{2},\ldots$. As long as the states that concern us are confined to small values of the quadrature variables these corrections can be ignored, but for states that extend out to larger values the corrections become significant and mean that $\hat{L}^{\mathrm{MB}}_{z}$ is modified from the expression given above to one which no longer corresponds to the canonical angular momentum operator.
@@
Fig. 2 plots the energies $\{\epsilon_{i}\}$ versus their $\langle\hat{L}_{z}^{\mathrm{MB}}\rangle$ values for a range of eigenstates of the Floquet operator [Eq. (2)]. Panel (a) is generated with no intra-species interactions ($u=0$) and shows the effect of having a finite number of particles in each BEC: the black dots are calculated numerically from the Floquet operator for $N=60$, whereas the red dots are $E_{0,m_{z}}$ with integer values of $m_{z}$ and correspond to the thermodynamic limit, $N\to\infty$, in which the effective Hamiltonian Eq. (10) becomes exact. One can see that for the lowest energy states $\langle\hat{L}_{z}^{\mathrm{MB}}\rangle$ gives values that agree well with integer values of $m_{z}$, but starting around $\langle\hat{L}_{z}^{\mathrm{MB}}\rangle=8$ the numerical and analytic results begin to disagree. In view of the previous discussion concerning finite size corrections these results are to be expected because states with lower angular momentum that have smaller orbits experience smaller finite size effects, whereas those with larger angular momentum and hence larger radial extent experience larger finite size effects.
@@
The disagreement is at its extreme in the top right of panel (a), where the energy required to increase the angular momentum by a small amount begins to diverge due to the hard wall of the 2D FSL. The approximate size of the eigenstates can be estimated from $\partial\psi_{0,m_{z}}/\partial r=0$ which gives $r_{\mathrm{size}}\approx\sqrt{\frac{m_{z}}{\omega}}$. For $u=0$, a rough upper bound on the angular momentum which the FSL can accommodate is given by the requirement that $r_{\mathrm{size}}<\sqrt{N/2}$ which means $m_{z}^{\mathrm{max}}=N\omega/2$. For the parameter values used in Fig. 2 we obtain $\langle\hat{L}_{z}^{\mathrm{MB}}\rangle_{\mathrm{max}}\approx 28.6$ which is below the upper bound of $m_{z}^{\mathrm{max}}=33$. Panel (b) shows the lowest energy states for different values of $u$. When for $u<1$ the ground state has zero angular momentum. However, when $u>1$ a dip can be seen at $\langle\hat{L}_{z}^{\mathrm{MB}}\rangle\approx 1$ indicating a ground state transition as $u$ is changed. Because the transition to nonzero angular momentum depends on the finite size of the system, it is not a quantum phase transition in the strict sense that requires the thermodynamic limit be taken.
@@
Figure 2: Floquet operator energies $\epsilon_{i}$ as a function of the expectation value of the many-body angular momentum operator $\langle\hat{L}_{z}^{\mathrm{MB}}\rangle$. The black dots in panel (a) give the five hundred lowest lying energy states found numerically for $N=60$, whereas the red dots are given by $E_{0,m_{z}}$ from Eq. (12) which corresponds to the case of an harmonic trap with no higher order trapping terms, i.e. assumes $N\rightarrow\infty$. Panel (b) shows a few of the lowest states for $N=60$ for different values of $u$ where in the case of $u=1.05$ the ground state develops nonzero angular momentum. In both panels $w=1$ and $\delta t=0.01$.
@@ IV.2 Ground state vortex transition
Figure 3: Density and phase plots of the ground state of the Floquet operator. Each row corresponds to a different value of $u$: $u=1.015$ for (a) and (b), $u=1.020$ for (c) and (d) and $u=1.035$ for (e) and (f). In the phase plots (right column), as $u$ increases, the ground state transitions from no central vortex in (b) to a singly quantized one in (d), then to a doubly quantized one in (f). Vortex cores are points of undefined phase, so the wave function must vanish at these points which results in the holes in the probability distributions in (c) and (e). The other parameters for the plots are $w=1$, $N=90$ and $\delta t=0.01$.
@@
Figure 4: Finite size scaling of the vortex transitions. The black curve in panel (a) plots the ground state expectation value of the many-body angular momentum operator $\langle\hat{L}^{\mathrm{MB}}_{z}\rangle_{\mathrm{GS}}$ versus $u$ which characterizes the intra-particle interaction strength, for $N=60$. The horizontal dashed red lines show the angular momenta values $m_{z}$ which are the quantum numbers relevant to the solutions of $\hat{H}_{\mathrm{eff}}$. Each step corresponds to an increase in the quantization of the central vortex. For larger $u$, the curve becomes smoother as the quantization breaks down due to finite size effects. (b) $|u_{i}-1|$ as a function of $1/N$ where $u_{i}$, $i=1,2,3$, are the values of $u$ where the first three steps occur and $u=1$ is the predicted vortex transition point. The curves are best fits of the data to polynomials of quadratic order. The parameter values are $w=1$ and $\delta t=0.01$.
@@
We now turn to studying the ground state of the Floquet operator in some detail. In Fig. 3 we plot both its probability distribution and its phase above and below the transition point. First, for $u<1$ in panel (a) we see that the ground state is simply a Gaussian centered at $m=n=0$ which agrees with the prediction in Eq. (13). In panel (b), the phase at the center shows no sign of a vortex, so this state has zero angular momentum. There are vortices near the edges of the FSL, however, the wave function is exponentially small in these regions, so they give negligible contribution to the angular momentum. By contrast, for $u>1$ panel (c) shows that the distribution has a hole at the center and the phase has also changed in (d), developing a singularity at the center where the synthetic magnetic field punctures the wave function. It is clearly shown that any circuit encircling the center once will accumulate a phase shift of $2\pi$, so this is a singly quantized vortex. Once more, the stability of this vortex state can be attributed to the existence of finite size effects. These effects lead to an increase in tunneling energy towards the system’s boundary resulting in confinement in Fock space that extends beyond harmonic trapping. As $u$ increases further, the ground state becomes more spread until its width is similar to the size of the doubly quantized vortex as seen in (e). The phase plot in (f) confirms that a path around the central phase singularity accumulates a $4\pi$ phase shift signalling the presence of two magnetic flux quanta.
@@
To further quantify the transition, we plot the ground state value of $\langle\hat{L}_{z}^{\mathrm{MB}}\rangle$ as a function of $u$ in Fig. 4 (a). As $u$ is increased, the ground state wave function undergoes a progressive widening in Fock space, such that as the width of the wave function approaches the size of the lowest angular momentum state there is sudden quantized jump in the value of $\langle\hat{L}_{z}^{\mathrm{MB}}\rangle$. With further increments in $u$, the wave function continues to spread, approaching the size of the second lowest angular momentum state, leading to a second jump, and so on. The red dashed horizontal lines are the integer values of the angular momentum predicted by $\hat{H}_{\mathrm{eff}}$ in Eq. (10). The agreement between the analytic and numerical values is best for low angular momenta, like the energies in Fig. 2 because these states are away from the edges. However, as $u$ increases eventually the width of the ground state approaches the size of the FSL and the quantization of the angular momentum breaks down as can be seen in the upper parts of the curve in Fig. 4 (a). It is worth noting that finite size effects are also responsible for the first jump not occurring precisely at the transition value of $u=1$.
@@ IV.3 Entanglement entropy between the two BECs
Quantized vortices in ordinary real space in BECs can often be accurately described within the Gross-Pitaevskii mean-field theory which does not involve entanglement. It is therefore interesting to ask whether this is also the case for Fock space vortices or whether there is intrinsically more entanglement that is generated. On the one hand, the effective Hamiltonian in Fock space given in Eq. (10) appears to be a single particle-type Hamiltonian. On the other hand, it should be remembered that the two Fock space coordinates $(x,y)$ each refer to different BECs and since vortices couple the two coordinates it seems that Fock space vortices may indeed be associated with entanglement between the two underlying species of atoms. We shall therefore consider the entanglement entropy between the two BECs as another possible measure to explore in relation to the vortex transition.
Figure 5: Ground state linear entanglement entropy between the two BECs as a function of the attractive intra-species interaction strength $u$. The steps in the linear entropy correspond to successive transitions to higher quantized vortex states. These steps occur at the same values of $u$ as the steps seen in Fig. 4. The horizontal dashed red lines are the predicted linear entanglement entropy values from Eq. (16). The parameter values are $w=1$, $N=60$ and $\delta t=0.01$.
For $u<1$, when there is no central vortex present, the ground state wave function takes the form of a 2D Gaussian located at the center of the FSL. This state is similar to that which would occur if the two BECs were not subject to driving or interacting with each other at all. Consequently, in this state, we anticipate no entanglement between the two BECs. For $u>1$ the two BECs must cooperate to form the vortex state and so we expect there to be some entanglement. According to Eq. (13), when $n_{r}=0$, the density matrix for any eigenstate of the system is $\rho=|m_{z}\rangle\langle m_{z}|$. In general, when one BEC is traced out, the resulting reduced density matrix, let it be $\rho_{x}$ corresponding only to the first BEC, will be mixed. The mixing can be quantified in terms of the linear entropy which is related to the purity of the reduced density matrix
@@ V Conclusion and Discussion
We have shown that the periodic driving of weak interactions between two species of BECs in a double well potential generates a weak synthetic magnetic field in the 2D FSL of the system. Like the effect of a magnetic field on electrons in a material, the synthetic magnetic field generates vortices in the wave function in the FSL leading to states with quantized angular momentum. With no or repulsive intra-species interactions, the ground state has zero angular momentum because the inherent confining potential in the FSL prevents the ground state wave function from spreading to a point where it can acquire one quantum of angular momentum. However, when the interactions are attractive and larger than $u=1$ it becomes energetically favorable for the ground state to spread and acquire nonzero angular momentum. This state is only stable for a finite number of particles in each BEC. In the thermodynamic limit, its angular momentum diverges because higher-order terms in the confining potential, which are needed to prevent the spreading of the wave function to infinity, vanish in this limit.
We note that the transition to a vortex state in Fock space under increasing interparticle interaction strength corresponds to a sudden jump in the occupation number difference between the original real space wells of the double well potential and is related to the macroscopic quantum self trapping transition that is known to occur in bosonic Josephson junctions [64, 65]. Here it is the version that occurs for attractive interactions and hence can occur in the ground state [69, 82, 83].
@@
A natural follow-on question to ask is whether interactions can occur in Fock space in order to generate an analogous quantity to a healing length and hence vortex lattices. Although there may be many ways to simulate interactions, the most obvious approach within the context of this paper involves introducing a second pair of two-mode BECs to serve as the degrees of freedom of a second particle in the FSL. Contact interactions between each pair of BECs can simulate interactions between the two FSL particles and periodic driving within each pair ensures both particles are influenced by a synthetic magnetic field. This approach can potentially lead to strongly correlated states (in Fock space) and hence pave the way for simulating more exotic forms of matter such as fractional quantum Hall states [85].

### v2

@@ IV.1 Eigenenergies and Eigenstates
In this paper we focus on the ground-state properties of the system, so we take $n_{r}=0$ from now on. When $u<1$, the ground-state energy is $E_{0,0}=\omega$ and the wave function is simply a Gaussian centered at $r=0$. In terms of the two BECs, it is centered at $m=n=0$ which is the state with an equal number of bosons of each BEC in each well. When $u>1$, the centrifugal limit is surpassed and Eq. (12) predicts that $E_{0,m_{z}}\to-\infty$ as $m_{z}\to\infty$. However, this is only true in the thermodynamic limit. For finite-size BECs there can be a unique ground state with $m_{z}\neq 0$ which, according to Eq. (13), is a vortex state with $m_{z}$ quanta of angular momentum.
@@
$\hat{L}^{\mathrm{MB}}_{z}$ represents the angular momentum operator in the FSL and we expect its mean value $\langle\hat{L}^{\mathrm{MB}}_{z}\rangle$ to accurately give $m_{z}$ when computed for the low-lying energy states well away from the edge of the FSL, which are the states of primary interest in this paper. However, we note that the correspondence between $\langle\hat{L}^{\mathrm{MB}}_{z}\rangle$ and $m_{z}$ can break down for higher states due to the omission of terms that depend on the number of particles as $1/N,1/N^{3/2},\ldots$. These terms were dropped in the derivation of $\hat{H}_{\mathrm{eff}}$ and lead to corrections that depend on the quadrature variables as $x^{2}y^{2},x^{2}p_{y}^{2},y^{2}p_{x}^{2},\ldots$. As long as the states that concern us are confined to small values of the quadrature variables these corrections can be ignored, but for states that extend out to larger values the corrections become significant and mean that $\hat{L}^{\mathrm{MB}}_{z}$ is modified from the expression given above to one which no longer corresponds to the canonical angular momentum operator.
@@
Fig. 2 plots the energies $\{\epsilon_{i}\}$ versus their $\langle\hat{L}_{z}^{\mathrm{MB}}\rangle$ values for a range of eigenstates of the Floquet operator [Eq. (2)]. Panel (a) is generated with no intra-species interactions ($u=0$) and shows the effect of having a finite number of particles in each BEC: the black dots are calculated numerically from the Floquet operator for $N=60$, whereas the red dots are $E_{0,m_{z}}$ with integer values of $m_{z}$ and correspond to the thermodynamic limit, $N\to\infty$, in which the effective Hamiltonian Eq. (10) becomes exact. One can see that for the lowest-energy states $\langle\hat{L}_{z}^{\mathrm{MB}}\rangle$ gives values that agree well with integer values of $m_{z}$, but starting around $\langle\hat{L}_{z}^{\mathrm{MB}}\rangle=8$ the numerical and analytic results begin to disagree. In view of the previous discussion concerning finite-size corrections these results are to be expected because states with lower angular momentum that have smaller orbits experience smaller finite-size effects, whereas those with larger angular momentum and hence larger radial extent experience larger finite size-effects.
@@
The disagreement is at its extreme in the top right of panel (a), where the energy required to increase the angular momentum by a small amount begins to diverge due to the hard wall of the 2D FSL. The approximate size of the eigenstates can be estimated from $\partial\psi_{0,m_{z}}/\partial r=0$ which gives $r_{\mathrm{size}}\approx\sqrt{\frac{m_{z}}{\omega}}$. For $u=0$, a rough upper bound on the angular momentum which the FSL can accommodate is given by the requirement that $r_{\mathrm{size}}<\sqrt{N/2}$ which means $m_{z}^{\mathrm{max}}=N\omega/2$. For the parameter values used in Fig. 2 we obtain $\langle\hat{L}_{z}^{\mathrm{MB}}\rangle_{\mathrm{max}}\approx 28.6$ which is below the upper bound of $m_{z}^{\mathrm{max}}=33$. Panel (b) shows the lowest-energy states for different values of $u$. When $u<1$ the ground state has zero angular momentum. However, when $u>1$ a dip can be seen at $\langle\hat{L}_{z}^{\mathrm{MB}}\rangle\approx 1$ indicating a ground-state transition as $u$ is changed. Because the transition to nonzero angular momentum depends on the finite size of the system, it is not a quantum phase transition in the strict sense that requires the thermodynamic limit be taken.
@@
Figure 2: Floquet operator energies $\epsilon_{i}$ as a function of the expectation value of the many-body angular momentum operator $\langle\hat{L}_{z}^{\mathrm{MB}}\rangle$. The black dots in panel (a) give the 500 lowest-lying energy states found numerically for $N=60$, whereas the red dots are given by $E_{0,m_{z}}$ from Eq. (12) which corresponds to the case of a harmonic trap with no higher-order trapping terms, i.e. assumes $N\rightarrow\infty$. Panel (b) shows a few of the lowest states for $N=60$ for different values of $u$ where in the case of $u=1.05$ the ground state develops nonzero angular momentum. In both panels $w=1$ and $\delta t=0.01$.
@@ IV.2 Ground state vortex transition
Figure 3: Density and phase plots of the ground state. Each row corresponds to a different value of $u$: $u=1.015$ for (a) and (b), $u=1.020$ for (c) and (d) and $u=1.035$ for (e) and (f). In the phase plots (right column), as $u$ increases, the ground state transitions from no central vortex in (b) to a singly quantized one in (d), then to a doubly quantized one in (f). Vortex cores are points of undefined phase, so the wave function must vanish at these points which results in the holes in the probability distributions in (c) and (e). The other parameters for the plots are $w=1$, $N=90$, and $\delta t=0.01$.
@@
Figure 4: Finite-size scaling of the vortex transitions. The black curve in panel (a) plots the ground-state expectation value of the many-body angular momentum operator $\langle\hat{L}^{\mathrm{MB}}_{z}\rangle_{\mathrm{GS}}$ vs $u$ which characterizes the intra-particle interaction strength, for $N=60$. The horizontal dashed red lines show the angular momenta values $m_{z}$ which are the quantum numbers relevant to the solutions of $\hat{H}_{\mathrm{eff}}$. Each step corresponds to an increase in the quantization of the central vortex. For larger $u$, the curve becomes smoother as the quantization breaks down due to finite-size effects. (b) $|u_{i}-1|$ as a function of $1/N$ where $u_{i}$, $i=1,2,3$, are the values of $u$ where the first three steps occur and $u=1$ is the predicted vortex transition point. The curves are best fits of the data to polynomials of quadratic order. The parameter values are $w=1$ and $\delta t=0.01$.
@@
We now turn to studying the ground state of the Floquet operator in some detail. In Fig. 3 we plot both its probability distribution and its phase above and below the transition point. First, for $u=1.015$ in panel (a) we see that the ground state is simply a Gaussian centered at $m=n=0$. In panel (b), the phase at the center shows no sign of a vortex, so this state has zero angular momentum. There are vortices near the edges of the FSL, however, the wave function is exponentially small in these regions, so they give negligible contribution to the angular momentum. We note that for finite sizes the transition point is shifted away from $u=1$ to a larger value which is why we get no vortex in panels (a) and (b) even though $u>1$. When $u=1.020$ panel (c) shows that the distribution has a hole at the center and the phase has also changed in (d), developing a singularity at the center where the synthetic magnetic field punctures the wave function. It is clearly shown that any circuit encircling the center once will accumulate a phase shift of $2\pi$, so this is a singly quantized vortex. Once more, the stability of this vortex state can be attributed to the existence of finite-size effects. These effects lead to an increase in tunneling energy towards the system’s boundary resulting in confinement in Fock space that extends beyond harmonic trapping. Finally, when $u=1.035$ the ground state becomes more spread until its width is similar to the size of the doubly quantized vortex as seen in (e). The phase plot in (f) confirms that a path around the central phase singularity accumulates a $4\pi$ phase shift signaling the presence of two magnetic flux quanta.
@@
To further quantify the transition, we plot the ground-state value of $\langle\hat{L}_{z}^{\mathrm{MB}}\rangle$ as a function of $u$ in Fig. 4 (a). As $u$ is increased, the ground-state wave function undergoes a progressive widening in Fock space, such that as the width of the wave function approaches the size of the lowest angular momentum state there is a sudden quantized jump in the value of $\langle\hat{L}_{z}^{\mathrm{MB}}\rangle$. With further increments in $u$, the wave function continues to spread, approaching the size of the second lowest angular momentum state, leading to a second jump, and so on. The red dashed horizontal lines are the integer values of the angular momentum predicted by $\hat{H}_{\mathrm{eff}}$ in Eq. (10). The agreement between the analytic and numerical values is best for low angular momenta, like the energies in Fig. 2 because these states are away from the edges. However, as $u$ increases, eventually the width of the ground state approaches the size of the FSL and the quantization of the angular momentum breaks down as can be seen in the upper parts of the curve in Fig. 4 (a). The connection between $u$ and the finite-size effects of the ground state comes from the fact that as $u$ increases the intra-species interactions become more attractive which promotes each BEC clumping into one well over the other. Having a large number of bosons in one well over the other corresponds to large values of $|m|$ and $|n|$, thus the wave function is closer to the edge of the FSL and therefore experiences larger finite-size effects. In the extreme example of taking the limit $u\to\infty$, we get $\hat{H}_{\mathrm{eff}}\to-\frac{u}{N}\left(\hat{J}_{z}^{2}+\hat{S}_{z}^{2}\right)$ which has degenerate ground states consisting of Fock states with $|m|=|n|=N/2$. They will have maximum disagreement with the eigenstates of the continuum Hamiltonian in Eq. (10) which has no concept of finite size.
@@ IV.3 Entanglement entropy between the two BECs
Quantized vortices in ordinary real space in BECs can often be accurately described within the Gross-Pitaevskii mean-field theory which does not involve entanglement. It is therefore interesting to ask whether this is also the case for Fock-space vortices or whether there is intrinsically more entanglement that is generated. On the one hand, the effective Hamiltonian in Fock space given in Eq. (10) appears to be a single particle-type Hamiltonian. On the other hand, it should be remembered that the two Fock-space coordinates $(x,y)$ each refer to different BECs and since vortices couple the two coordinates it seems that Fock-space vortices may indeed be associated with entanglement between the two underlying species of atoms. We shall therefore consider the entanglement entropy between the two BECs as another possible measure to explore in relation to the vortex transition.
Figure 5: Ground-state linear entanglement entropy between the two BECs as a function of the attractive intra-species interaction strength $u$. The steps in the linear entropy correspond to successive transitions to higher quantized vortex states. These steps occur at the same values of $u$ as the steps seen in Fig. 4. The horizontal dashed red lines are the predicted linear entanglement entropy values from Eq. (16). The parameter values are $w=1$, $N=60$ and $\delta t=0.01$.
For $u<1$, when there is no central vortex present, the ground-state wave function takes the form of a 2D Gaussian located at the center of the FSL. This state is similar to that which would occur if the two BECs were not subject to driving or interacting with each other at all. Consequently, in this state, we anticipate no entanglement between the two BECs. For $u>1$ the two BECs must cooperate to form the vortex state and so we expect there to be some entanglement. According to Eq. (13), when $n_{r}=0$, the density matrix for any eigenstate of the system is $\rho=|m_{z}\rangle\langle m_{z}|$. In general, when one BEC is traced out, the resulting reduced density matrix, let it be $\rho_{x}$ corresponding only to the first BEC, will be mixed. The mixing can be quantified in terms of the linear entropy which is related to the purity of the reduced density matrix
@@ V Conclusion and Discussion
We have shown that the periodic driving of weak interactions between two species of BECs in a double-well potential generates a weak synthetic magnetic field in the 2D FSL of the system. Like the effect of a magnetic field on electrons in a material, the synthetic magnetic field generates vortices in the wave function in the FSL leading to states with quantized angular momentum. With no or repulsive intra-species interactions, the ground state has zero angular momentum because the inherent confining potential in the FSL prevents the ground-state wave function from spreading to a point where it can acquire one quantum of angular momentum. However, when the interactions are attractive and larger than $u=1$ it becomes energetically favorable for the ground state to spread and acquire nonzero angular momentum. This state is only stable for a finite number of particles in each BEC. In the thermodynamic limit, its angular momentum diverges because higher-order terms in the confining potential, which are needed to prevent the spreading of the wave function to infinity, vanish in this limit.
We note that the transition to a vortex state in Fock space under increasing interparticle interaction strength corresponds to a sudden jump in the occupation number difference between the original real-space wells of the double-well potential and is related to the macroscopic quantum self trapping transition that is known to occur in bosonic Josephson junctions [64, 65]. Here it is the version that occurs for attractive interactions and hence can occur in the ground state [69, 93, 94].
@@
A natural follow-up question to ask is whether interactions can occur in Fock space in order to generate an analogous quantity to a healing length and hence vortex lattices. Although there may be many ways to simulate interactions, the most obvious approach within the context of this paper involves introducing a second pair of two-mode BECs to serve as the degrees of freedom of a second particle in the FSL. Contact interactions between each pair of BECs can simulate interactions between the two FSL particles and periodic driving within each pair ensures both particles are influenced by a synthetic magnetic field. This approach can potentially lead to strongly correlated states (in Fock space) and hence pave the way for simulating more exotic forms of matter such as fractional quantum Hall states [96].